Book Co. has 1.0 million shares of common equity with a par​ (book) value of $1.00​, retained earnings of $30.0 million, and its shares have a market value of $50.00 per share. It also has debt with a par value of $20.0 million that is trading at 101% of par.

a. What is the market value of its​ equity?

b. What is the market value of its​ debt?

c. What weights should it use in computing its​ WACC?

a. What is the market value of its​ equity?

The market value of the equity is ____? million. ​(Round to two decimal​ places.)
Answer #1

a.

Market Value of Equity = Number of Shares(Price per Share)

Market Value of Equity = 1,000,000(50)

Market Value of Equity = $50,000,000

b.

Market Value of Debt = 20,000,000(1.01)

Market Value of Debt = $20,200,000

c.

Weight of Equity = 50,000,000/(50,000,000 + 20,200,000)

Weight of Equity = 71.23%

Weight of Debt = 20,200,000/(50,000,000 + 20,200,000)

Weight of Debt = 28.77%